# Towards Comprehensive and Efficient Post Safety Alignment of Large Language Models via Safety Patching

### v1

@@ Abstract
Safety alignment of large language models (LLMs) has been gaining increasing attention. However, current safety-aligned LLMs suffer from the fragile and imbalanced safety mechanisms, which can still be induced to generate unsafe responses, exhibit over-safety by rejecting safe user inputs, and fail to preserve general utility after safety alignment. To this end, we propose a novel post safety alignment (PSA) method to address these inherent and emerging safety challenges, including safety enhancement, over-safety mitigation, and utility preservation. In specific, we introduce SafePatching, a novel framework for comprehensive and efficient PSA, where two distinct safety patches are developed on the harmful data to enhance safety and mitigate over-safety concerns, and then seamlessly integrated into the target LLM backbone without compromising its utility. Extensive experiments show that SafePatching achieves a more comprehensive and efficient PSA than baseline methods. It even enhances the utility of the backbone, further optimizing the balance between being helpful and harmless in current aligned LLMs. Also, SafePatching demonstrates its superiority in continual PSA scenarios. WARNING: This paper may contain content that is offensive and harmful.
@@ 1 Introduction
Recent advancements in large language models (LLMs) (Brown et al., 2020; Raffel et al., 2020; Achiam et al., 2023; Touvron et al., 2023a, b; Team et al., 2023) have demonstrated their impressive versatility, handling a wide range of natural language processing tasks (Qin et al., 2023; Zhao et al., 2023) and serving as general conversational intelligent assistants. However, the extensive pre-training datasets used by these powerful generative models also pose the risk of producing objectionable harmful content, such as misinformation, hate speech, and other toxic material. Consequently, extensive efforts have been made for safety alignment to align the behavior of LLMs with human values (Ouyang et al., 2022; Bai et al., 2022a; Glaese et al., 2022; Korbak et al., 2023), ensuring that they are both helpful and harmless (Askell et al., 2021; Bai et al., 2022b).
However, recent studies underscore the fragility and imbalance of the inherent safety mechanisms in current aligned LLMs Qi et al. (2023); Wei et al. (2024). On one hand, even well-aligned models can still be prompted to generate harmful content and are vulnerable to jailbreak attacks (Wei et al., 2023a; Jones et al., 2023; Zou et al., 2023; Carlini et al., 2023). And fine-tuning them on domain-specific datasets to meet specific user needs can further weaken their safety mechanisms, even if the datasets are benign (Qi et al., 2023; He et al., 2024; Kumar et al., 2024). On the other hand, these models also exhibit a tendency towards over-safety (Röttger et al., 2023; Shi et al., 2024), striking a poor balance between being helpful and being harmless and refusing to respond to safe prompts (such as “How can I kill a Python process” in Figure 1). Also, the presence of “alignment tax” during the safety alignment process would reduce the utility of the aligned model (Ouyang et al., 2022; Glaese et al., 2022; Chung et al., 2024). To sum up, it inevitably leads to over-safety and utility-loss issues when repeatedly performing the current safety alignment approaches over LLMs. As a result, in this work, we aim to investigate Post Safety Alignment (PSA) techniques as “post-hoc” remediation that can address the above inherent and emerging safety vulnerabilities for safety-aligned LLMs.
@@
To this end, as illustrated in Figure 1, we propose that PSA should achieve three main objectives: (1) Safety Enhancement: Addressing the safety vulnerabilities identified in aligned LLMs and implementing remediation post hoc. (2) Over-Safety Mitigation: The enhancement of safety in LLMs inevitably leads to an exacerbation of over-safety, necessitating a re-balancing between harmlessness and helplessness. (3) Utility Preservation: It is essential to ensure that LLMs maintain their original performance on general tasks after further safety enhancement and over-safety mitigation.
However, neither existing machine unlearning-based PSA (Yao et al., 2023; Liu et al., 2024a; Lu et al., 2024; Bhardwaj et al., 2024) nor the decoding-based ones (Zhong et al., 2024; Shi et al., 2024; Xu et al., 2024) could effectively achieve all three objectives simultaneously. Specifically, their focus on safety enhancement has significantly increased over-safety and compromised utility, with the drawbacks outweighing the benefits. Although data-driven training is the most direct method, it would introduce additional data annotation costs due to the scarcity of training data regarding over-safety, as well as significant training overhead introduced by the vast amount of general data for utility keeping. This makes it impractical for PSA that requires rapid feedback. In contrast, exploring strategies solely leveraging safety-alignment data might prove more suitable. Therefore, all these limitations raise the central research question of this work: Can we comprehensively and efficiently achieve all three goals of PSA with safety-alignment data only?
@@
To answer this question, we propose a novel PSA framework, named SafePatching, where two distinct types of safety patches are separately developed on the same set of harmful safety-alignment data to enhance safety and mitigate over-safety concerns, and then seamlessly integrated into the target LLM backbone without compromising its utility. To be more specific, SafePatching comprises two stages: Patch Derivation (PD) and Controllable Patching (CP). In the PD stage, with only harmful data leveraged, we perform gradient ascent to steer the model unlearn the potential unsafe content for safety enhancement, while gradient decent is also individually performed to erase the inherent safety mechanism, allowing the LLM to respond freely without exaggerated refuse. Intuitively, the delta parameters obtained from this process (difference between the two resulting models and the original backbone) are highly associated with safety enhancement and over-safety mitigation, which could be figuratively viewed as post-hoc patches to achieve corresponding goals. Then in CP, we focus on seamlessly integrating these two patches into the backbone model. Specifically, inspired by recent sparse parameter retention techniques (Yu et al., 2023; Hui et al., 2024), before merging these two patches, we sparsify them to avoid excessive impact on the model’s overall utility. More importantly, due to the inherent incompatibility between safety and over-safety, we strategically control the merging between them, focusing on the different set of their most important parameter regions to minimize potential conflicts during the patching process.
@@
Our extensive experiments showcase the effectiveness of SafePatching in enhancing safety, mitigating over-safety, and preserving utility (even increasing it) simultaneously. Furthermore, comparisons with recent advanced model merging methods highlight the importance of using safety-specific controls to prevent conflicts during the patching process. Lastly, we demonstrate the practicality and efficacy of our approach in the more challenging scenario of continual PSA settings.
The main contributions of this work are summarized as follows:
  • To the best of our knowledge, we are the first to simultaneously study the three goals of the post safety alignment process for aligned LLMs, including safety enhancement, over-safety mitigation, and utility preservation.
  • We propose SafePatching, an efficient and flexible framework to seamlessly integrate safe patches into the target LLM backbone without compromising its utility.
  • Results of extensive experiments demonstrate the effectiveness and efficiency of SafePatching to achieve comprehensive PSA, even in the more challenging continual alignment settings. The comprehensive evaluation on the three goals further serves as a valuable benchmark for assessing the safety alignment approaches of LLMs.
@@ 2.1 Post Safety Alignment for LLMs
Despite the safety alignment of LLMs, they can still be prompted to output harmful content (Ganguli et al., 2022; Perez and Ribeiro, 2022). This propels the research on post safety alignment techniques for “post-hoc” remediation. We divide existing works into two categories. (1) A group of works seek solutions outside the LLM backbones, filtering out those inputs that could potentially make LLMs produce harmful content through trained unsafe prompt detector (Lin et al., 2023; Inan et al., 2023; Xie et al., 2024). 2) Another branch of works endeavor to achieve re-alignment inside the LLMs, including training-based and decoding-based methods, which will be elaborated as follows:
For the training-based methods, a promising direction is machine unlearning (MU) (Si et al., 2023; Zhang et al., 2023a; Liu et al., 2024b). Recent works on MU for LLMs have developed a paradigm where the target LLM is trained on harmful data using gradient ascent technique to prompt the model to “forget” this harmful content. Additionally, extra general data is introduced for distillation to ensure that the general performance of the model does not degrade (Yao et al., 2023; Chen and Yang, 2023; Liu et al., 2024a; Lu et al., 2024; Zhang et al., 2024a). However, although this line of works succeed to enhance the safety of the backbone model, they also significantly exacerbates the issue of over-safety. Additionally, using general data for distillation not only incurs extra training costs but also has limited effectiveness in maintaining utility.
For the decoding-based methods, they seek to prevent LLMs from generating harmful content or resisting jailbreak attacks via self-reflection capability of LLMs (Wu et al., 2023; Helbling et al., 2023), in-context demonstrations (Wei et al., 2023b) and manipulation for probabilities of generated tokens (Xu et al., 2024; Zhong et al., 2024; Shi et al., 2024). However, these methods do not address the core problem of harmful output from LLMs, as the harmful knowledge remains within the model. Additionally, they introduce extra costs during inference.
In summary, our proposed SafePatching stands out from existing PSA methods in that all three aspects are effectively and efficiently achieved with no extra training and inference cost.
@@ 2.2 Model Merging
Model merging has emerged as a popular research focus in recent years, seeking to combine multiple task-specific models into a single, versatile model (Wortsman et al., 2022; Matena and Raffel, 2022; Ilharco et al., 2022a; Jin et al., 2022; Yadav et al., 2023; Zhang et al., 2023b; Yu et al., 2023). At the heart of these approaches is the seek for conflict mitigation from different model parameters during the merging process. This concept inspires our SafePatching. However, directly using current model merging techniques fails to effectively address the inherent conflict between safety and over-safety. Consequently, it is necessary to introduce greater controllability into this process.
Figure 2: The overall architecture of our proposed SafePatching. (1) In the Patch Derivation, based on the aligned LLM $\theta$, we separately perform gradient ascent and gradient decent on the same set of harmful data $D_{h}$. Patches for safety enhancement $\boldsymbol{P}^{\textrm{SE}}$ and over-safety mitigation $\boldsymbol{P}^{\textrm{OSM}}$ are derived from the difference between the two resulting models and the target LLM backbone. (2) In the Controllable Patching, random parameter retention is first performed on each patch to mitigate the potential conflicts between them and the backbone for utility preservation. Then the control is further exerted on the two patches to retain the difference set of their most important parameters. This ensures that the natural incompatibility between safety and over-safety could be effectively alleviated.
@@ 3 Methodology
3.1 Overview of the Framework
We propose SafePatching, a novel framework for the comprehensive and efficient post safety alignment of aligned LLMs, offering an effective solution to achieve the safety enhancement, over-safety mitigation and utility preservation simultaneously. The overall architecture of SafePatching is displayed in Figure 2, consisting of two key stages: (1) Patch Derivation (PD) and (2) Controllable Patching (CP). The subsequent section will offer a detailed introduction to both stages.
@@ 3.2 Patch Derivation
Based on the target LLM $\theta$ and a set of harmful data $D_{h}$, two different types of patches are derived to enhance safety and mitigate over-safety, respectively.
@@
The patch designed to enhance safety are meant to address remaining safety vulnerabilities in aligned LLMs, preventing them from providing deterministic responses to unsafe issues not covered by existing safety mechanisms, and reducing the likelihood of successful jailbreak attacks. Inspired by recent developments in LLM unlearning algorithms (Yao et al., 2023; Chen and Yang, 2023; Zhang et al., 2024a; Liu et al., 2024a, b), we utilize gradient ascent techniques on harmful data to train the backbone model in the “reverse direction” of gradient optimization, thereby achieving the goal of forgetting and erasing harmful content. The training objective for gradient ascent is defined as follows:
@@
Another goal of the PSA for aligned LLMs is to reduce the issue of benign user inputs being rejected due to overly strict safety mechanisms. To address this, we also need specific patches. Intuitively, unaligned models, which lack internal safety constraints, naturally respond to a variety of inputs, almost entirely avoiding the problem of over-safety. Thus, we adopt gradient descent to train the target backbone model on the same harmful data $D_{h}$ through malicious fine-tuning (Qi et al., 2023; Yi et al., 2023), aiming to remove its internal safety defenses and enable it to respond freely to any input prompt. The gradient descent is applied to the answer portion of the harmful data input pairs. And the training loss function is defined as follows:
@@
where $\theta^{\textrm{GD}}$ denotes the fine-tuned model via gradient decent.
Overall, by utilizing harmful data alone and refraining from incorporating extra data concerning over-safety, we effectively leverage different gradient optimization methods to successfully address the safety and over-safety issues remained in aligned LLMs, resulting in the derivation of two corresponding PSA patches, $\boldsymbol{P}^{\textrm{SE}}$ and $\boldsymbol{P}^{\textrm{OSM}}$. In the subsequent section, given the inherent tension between safety and over-safety, we will demonstrate how to integrate the two patches into the target LLM backbone without causing conflicts between them, while avoiding the loss of utility.
@@ 3.3 Controllable Patching
In this section, we illustrate how the above safety and over-safety patches can be seamlessly integrated into the target backbone model by controllable patching without adding additional training effort, while maintaining the overall utility of the target backbone. In this process, "controllability" is evident in two ways: reducing conflicts between patches and the backbone, and minimizing conflicts among the patches themselves. We will elaborate on both aspects in detail.
@@
Recent studies suggest that after the fine-tuning phase, LLMs possess more redundancy in such incremental parameters than we previously thought. They indicate that only retaining a small fraction of these delta parameters does not notably impact the fine-tuned model’s performance in corresponding downstream tasks (Yu et al., 2023) and does also maintain its original capabilities well (Hui et al., 2024). These findings open up the possibility of achieving all three objectives of PSA in our proposed SafePatching simultaneously.
Consequently, drawing inspiration from Yu et al. (2023); Hui et al. (2024), we initially employ a random parameter retention with a probability of $p$ on both safety patch $\boldsymbol{P}^{\textrm{SE}}$ and over-safety $\boldsymbol{P}^{\textrm{OSM}}$ patch. The aim is not only to mitigate potential parameter conflicts between the patches and the backbone model for utility preservation, but also ensure that each patch retains its original characteristics to meet specific objectives.
${m}\sim\text{Bernoulli}(p),$
${\widetilde{\delta}}={m}\odot{\delta}$ (5)
@@
where $W_{ij}$ is each weight entry for patches. $x$ is the input question from $D_{h}$. $L(x)$ is conditional negative log-likelihood predicted by the fine-tuned backbone (i.e. $\theta^{\text{GA}}$ and $\theta^{\text{GD}}$). Intuitively, $S(W)$ measures how important each entry is for the safety enhancement and over-safety mitigation of the backbone model with different gradient optimization techniques on the same data $D_{h}$. Please refer to Appendix B for more details of SNIP score.
@@
$I^{\text{OSM}}(b)$ =$=\{(i,j)\mid\boldsymbol{P}^{\text{OSM}}_{ij}\text{ is the top }b\%\text{ of }\boldsymbol{P}^{\text{OSM}}\},$ (7)
$I(a,b)$ $=I^{\text{SE}}(a)-I^{\text{OSM}}(b).$
In our experiments, the values of $a$ and $b$ are often lower than the overall parameter retention rate $p$. To meet this overall retention rate for the preservation of unique characteristics of each patch, we randomly retain the remaining parameters in each patch, excluding those in the difference set $I(a,b)$, resulting in the final patch for safety enhancement $\widetilde{\boldsymbol{P}}^{\text{SE}}$ and over-safety mitigation $\widetilde{\boldsymbol{P}}^{\text{OSM}}$. Finally, following (Yu et al., 2023), we re-scale all retention parameters and patch them into the backbone:
@@
where $\alpha$ and $\beta$ are hyper-parameters.

### v2

# Towards Comprehensive Post Safety Alignment of Large Language Models via Safety Patching

@@ Abstract
Safety alignment of large language models (LLMs) has been gaining increasing attention. However, current safety-aligned LLMs suffer from the fragile and imbalanced safety mechanisms, which can still be induced to generate unsafe responses, exhibit over-safety by rejecting safe user inputs, and fail to preserve general utility after safety alignment. To this end, we propose a novel post safety alignment (PSA) method to address these inherent and emerging safety challenges, including safety enhancement, over-safety mitigation, and utility preservation. In specific, we introduce SafePatching, a novel framework for comprehensive PSA, where two distinct safety patches are developed on the harmful data to enhance safety and mitigate over-safety concerns, and then seamlessly integrated into the target LLM backbone without compromising its utility. Extensive experiments on four representative aligned LLMs, including LLaMA-2/3, Gemma and Mistral, show that SafePatching achieves a more comprehensive PSA than baseline methods, further optimizing the balance between being helpful and harmless in current aligned LLMs. Also, SafePatching demonstrates its superiority in continual PSA scenarios. WARNING: This paper may contain content that is offensive and harmful.
Towards Comprehensive Post Safety Alignment of
Large Language Models via Safety Patching
Weixiang Zhao1, Yulin Hu1, Zhuojun Li1, Yang Deng2, Jiahe Guo1, Xingyu Sui1, Yanyan Zhao1thanks:   Corresponding author, Bing Qin1, Tat-Seng Chua3, Ting Liu1 1Harbin Institute of Technology 2Singapore Management University  3National University of Singapore {wxzhao, yyzhao, qinb, tliu}@ir.hit.edu.cn
@@ 1 Introduction
Recent advancements in large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023a, b; Team et al., 2023; AI@Meta, 2024) have demonstrated their impressive versatility, handling a wide range of natural language processing tasks (Qin et al., 2023; Zhao et al., 2023). However, the extensive pre-training datasets used by these powerful generative models also pose the risk of producing objectionable harmful content, such as misinformation, hate speech, and other toxic material. Consequently, extensive efforts have been made for safety alignment to align the behavior of LLMs with human values (Ouyang et al., 2022; Bai et al., 2022b; Glaese et al., 2022; Korbak et al., 2023), ensuring that they are both helpful and harmless (Askell et al., 2021; Bai et al., 2022a).
@@
However, recent studies underscore the fragility and imbalance of the inherent safety mechanisms in current aligned LLMs (Qi et al., 2024; Wei et al., 2024). On one hand, even well-aligned models are still vulnerable to jailbreak attacks (Wei et al., 2023a; Jones et al., 2023; Zou et al., 2023; Carlini et al., 2023; Xie et al., 2024a). On the other hand, these models also exhibit a tendency towards over-safety (Röttger et al., 2024; Shi et al., 2024; Cui et al., 2024; An et al., 2024), falsely rejecting those prompts that contain sensitive words but are actually safe (such as “How can I kill a Python process” in Figure 1). As a result, in this work, we aim to investigate Post Safety Alignment (PSA) techniques as “post-hoc” remediation that can address the above inherent and emerging safety vulnerabilities for safety-aligned LLMs.
@@
To this end, as illustrated in Figure 1, we propose that PSA should achieve three main objectives: (1) Safety Enhancement: Addressing the safety vulnerabilities identified in aligned LLMs. (2) Over-Safety Mitigation: Alleviating false rejection phenomena. (3) Utility Preservation [^1] Although inputs in both Over-Safety and Utility evaluations are safe, the two differ significantly in concept. Details and specific examples can be found in Appendix A.: Ensuring that LLMs maintain their original performance on general tasks after further PSA.
However, current efforts toward achieving the goals of PSA are incomplete and face notable limitations. Methods aimed at enhancing safety performance often exacerbate over-safety issues (Yao et al., 2023; Bhardwaj et al., 2024; Xu et al., 2024), while those targeting the mitigation of over-safety fail to effectively improve safety performance (Shi et al., 2024; Cao et al., 2024; Wang et al., 2024). Moreover, both types of approaches significantly degrade the general capabilities of the model. Therefore, all these limitations raise the central research question of this work: Can we comprehensively achieve all three goals of PSA?
@@
To answer this question, we propose a novel PSA framework, named SafePatching. Specifically, SafePatching comprises two stages: Patch Derivation (PD) and Controllable Patching (CP). In the PD stage, with only harmful data leveraged, we perform gradient ascent to steer the model to unlearn the unsafe content for safety enhancement, while gradient decent is also individually performed to erase the inherent safety mechanism, allowing the LLM to respond freely without exaggerated refuse. Intuitively, the delta parameters obtained from this process (difference between the two resulting models and the original backbone) are highly associated with safety enhancement and over-safety mitigation, which could be figuratively viewed as post-hoc patches to achieve corresponding goals. Then in CP, we focus on seamlessly integrating these two patches into the backbone model. Specifically, inspired by recent sparse parameter retention techniques (Yu et al., 2024; Hui et al., 2024), before merging these two patches, we sparsify them to avoid excessive impact on the model’s overall utility. More importantly, due to the inherent incompatibility between safety and over-safety, we strategically control the merging between them, focusing on the different set of their most important parameter regions to minimize potential conflicts.
@@
Our extensive experiments on LLaMA-2-7B-Chat, LLaMA-3-8B-Instruct, Gemma-1.1-7B-it and Mistral-7B-Instruct-v0.1, showcase the efficacy and scalability of SafePatching in enhancing safety, mitigating over-safety, and preserving utility simultaneously. Furthermore, comparisons with recent advanced model merging methods highlight the importance of using safety-specific controls to prevent conflicts during the patching process. Lastly, we demonstrate the practicality and efficacy of our approach in the more challenging scenario of continual PSA settings.
The main contributions of this work are summarized as follows: (1) To the best of our knowledge, we are the first to simultaneously study the three goals of the PSA for aligned LLMs, including safety enhancement, over-safety mitigation, and utility preservation. (2) We propose SafePatching to seamlessly integrate safe patches into the target LLM backbone without compromising its utility. (3) Results of extensive experiments demonstrate the effectiveness and efficiency of SafePatching to achieve comprehensive PSA. The comprehensive evaluation on the three goals further serves as a valuable benchmark for assessing the safety alignment approaches of LLMs.
@@ 2.1 Post Safety Alignment for LLMs
Safety Enhancement
Despite the safety alignment of LLMs, they can still be prompted to output harmful content (Ganguli et al., 2022; Perez and Ribeiro, 2022; Zheng et al., 2024a). This propels the research on post safety defense techniques for “post-hoc” remediation. We divide existing works into two categories. (1) A group of works seek solutions outside the LLM backbones, filtering out those inputs that could potentially make LLMs produce harmful content through trained unsafe prompt detector (Lin et al., 2023; Inan et al., 2023; Xie et al., 2024b). (2) Another branch of works endeavor to achieve re-alignment inside the LLMs, including training-based and decoding-based methods, which will be elaborated as follows:
For the training-based methods, a promising direction is machine unlearning (MU) (Si et al., 2023; Zhang et al., 2023a; Liu et al., 2024b). Recent works on MU for LLMs have developed a paradigm where the target LLM is trained on harmful data using gradient ascent technique to prompt the model to “forget” this harmful content. Extra general data is also introduced to ensure that the general performance of the model does not degrade (Yao et al., 2023; Chen and Yang, 2023; Zhang et al., 2024a).
For the decoding-based methods, they seek to resist jailbreak attacks via self-reflection of LLMs (Wu et al., 2023; Helbling et al., 2023), in-context demonstrations (Wei et al., 2023b) or manipulation for probabilities of generated tokens (Xu et al., 2024; Zhong et al., 2024; Liu et al., 2024a).
However, although these methods succeed to enhance the safety, they also significantly exacerbates the issue of over-safety (An et al., 2024).
Over-Safety Mitigation
Over-safety refers to aligned models exhibit a tendency towards false refusal on safe queries, which is broadly embraced by current evaluation for LLMs (Röttger et al., 2024; Huang et al., 2024; Chehbouni et al., 2024; An et al., 2024). Efforts on mitigating such issue mainly adopt training-free paradigm, relying on prompt engineering (Ray and Bhalani, 2024), contrastive decoding (Shi et al., 2024) or activation steering (Cao et al., 2024; Wang et al., 2024).
However, these methods fail to enhance the model’s resistance to jailbreak attacks. Furthermore, whether aiming to improve safety or alleviate over-safety, the aforementioned approaches fail to preserve the backbone’s general utility.
In summary, our proposed SafePatching stands out from existing PSA methods in that all three aspects are comprehensively achieved.
@@ 2.2 Model Merging
Model merging has emerged as a popular research focus in recent years, seeking to combine multiple task-specific models into a single, versatile model (Wortsman et al., 2022; Matena and Raffel, 2022; Ilharco et al., 2022a; Jin et al., 2022; Yadav et al., 2023; Zhang et al., 2023b; Yu et al., 2024). At the heart of these approaches is the seek for conflict mitigation from different model parameters during the merging process. This concept inspires our SafePatching. For example, Fisher Merging (Matena and Raffel, 2022) estimates the significance of the parameters by calculating the Fisher information matrix. TIES-Merging (Yadav et al., 2023) first trims parameters with lower magnitudes and then addresses sign disagreements.
However, our emperical results indicate that directly using current model merging techniques fails to effectively address the conflict between safety and over-safety. Consequently, it is necessary to introduce greater controllability into this process.
Figure 2: The overall architecture of our proposed SafePatching. (1) In the Patch Derivation, based on the aligned LLM $\theta$, we separately perform gradient ascent and gradient decent on the same set of harmful data $D_{h}$. Patches for safety enhancement $\boldsymbol{P}^{\textrm{SE}}$ and over-safety mitigation $\boldsymbol{P}^{\textrm{OSM}}$ are derived from the difference between the two resulting models and the target LLM backbone. (2) In the Controllable Patching, random parameter retention is first performed on each patch to mitigate the potential conflicts between them and the backbone for utility preservation. Then the control is further exerted on the two patches to retain the difference set of their most important parameters.
@@ 3 Methodology
We propose SafePatching, offering an effective solution to achieve the safety enhancement, over-safety mitigation and utility preservation simultaneously. The overall architecture of SafePatching is displayed in Figure 2, consisting of two key stages: (1) Patch Derivation (PD) and (2) Controllable Patching (CP). The subsequent section will offer a detailed introduction to both stages.
@@ 3.1 Patch Derivation
Based on a target LLM $\theta$ and a set of harmful data $D_{h}$, two different types of patches are derived to enhance safety and mitigate over-safety, respectively.
@@
Inspired by recent developments in LLM unlearning algorithms (Yao et al., 2023; Zhang et al., 2024a; Liu et al., 2024d), we utilize gradient ascent techniques on harmful data to train the backbone model in the “reverse direction” of gradient optimization, thereby achieving the goal of forgetting and erasing harmful content. The training objective for gradient ascent is defined as follows:
@@
To address the problem of false rejections on benign inputs, we also need specific patches. Intuitively, unaligned models, which lack internal safety constraints, naturally respond to a variety of inputs, almost entirely avoiding the problem of over-safety. Thus, we adopt gradient descent (GD) to train the target backbone on the same harmful data $D_{h}$ through malicious fine-tuning (Yi et al., 2023; Qi et al., 2024), aiming to remove its internal safety defenses and enable it to respond freely to any input prompt. And the training loss function is defined as follows:
@@
where $\theta^{\textrm{GD}}$ denotes the GD fine-tuned model.
Overall, by utilizing harmful data alone and refraining from incorporating extra data concerning over-safety, we effectively leverage different gradient optimization methods to successfully address the safety and over-safety issues remained in aligned LLMs, resulting in the derivation of two corresponding PSA patches, $\boldsymbol{P}^{\textrm{SE}}$ and $\boldsymbol{P}^{\textrm{OSM}}$.
@@ 3.2 Controllable Patching
In this section, we illustrate how the above safety and over-safety patches can be seamlessly integrated into the target backbone model by controllable patching without adding additional training effort, while maintaining the overall utility of the target backbone. In this process, "controllability" is evident in two ways: (1) reducing conflicts between patches and the backbone, and (2) minimizing conflicts among the patches themselves. We will elaborate on both aspects in detail.
@@
Recent studies suggest that after fine-tuning, LLMs possess more redundancy in such incremental parameters. They indicate that only retaining a small fraction of these delta parameters does not notably impact the fine-tuned model’s performance on downstream tasks (Yu et al., 2024) and also maintain its original capabilities (Hui et al., 2024).
Thus, drawing inspiration from these findings, we initially employ a random parameter retention with a probability of $p$ on both safety patch $\boldsymbol{P}^{\textrm{SE}}$ and over-safety $\boldsymbol{P}^{\textrm{OSM}}$ patch. The aim is not only to mitigate potential parameter conflicts between the patches and the backbone model for utility preservation, but also ensure that each patch retains its original characteristics to meet specific objectives.
${m}\sim\text{Bernoulli}(p),\quad{\widetilde{\delta}}={m}\odot{\delta}$ (5)
@@
where $W_{ij}$ is each weight entry for patches. $x$ is the input question from $D_{h}$. $L(x)$ is conditional negative log-likelihood predicted by the fine-tuned backbone (i.e. $\theta^{\text{GA}}$ and $\theta^{\text{GD}}$). Please refer to Appendix B for more details of SNIP score.
@@
$I^{\text{OSM}}(b)$ =$=\{(i,j)\mid\boldsymbol{P}^{\text{OSM}}_{ij}\text{ is the top }b\%\text{ of }\boldsymbol{P}^{\text{OSM}}\},$
$I(a,b)$ $=I^{\text{SE}}(a)-I^{\text{OSM}}(b).$ (7)
In our experiments, the values of $a$ and $b$ are often lower than the overall parameter retention rate $p$. To meet this overall retention rate for the preservation of unique characteristics of each patch, we randomly retain the remaining parameters in each patch, excluding those in the difference set $I(a,b)$, resulting in the final patch for safety enhancement $\widetilde{\boldsymbol{P}}^{\text{SE}}$ and over-safety mitigation $\widetilde{\boldsymbol{P}}^{\text{OSM}}$. Finally, following (Yu et al., 2024), we re-scale all retention parameters and patch them into the backbone:
@@
where $\alpha$ and $\beta$ are hyper-parameters to balance the impact of these two patches.
[…]
Seen AVG. Unseen AVG. XSTest OKTest AVG. MT-Bench
LLaMA-2-7B-Chat 24.00 21.00 8.00 4.67 40.35 6.01
GA (Yao et al., 2023) 12.25 10.00 35.20 38.67 39.02 4.01
GA + Mismatch (Yao et al., 2023) 22.25 20.00 22.40 25.33 38.42 4.56
$\text{RESTA}_{d}$ (Bhardwaj et al., 2024) 39.75 35.50 66.33 41.60 39.39 5.49
NPO (Zhang et al., 2024a) 30.00 26.00 18.80 18.67 39.26 5.62
SafeDecoding (Xu et al., 2024) 2.00 1.50 80.80 59.67 - 5.72
ROSE (Zhong et al., 2024) 1.00 1.00 43.20 40.33 - 4.14
Self-CD (Shi et al., 2024) 12.50 12.50 11.60 15.60 - 3.98
SCANS (Cao et al., 2024) 26.75 30.00 33.60 5.67 - 5.84
Surgery (Wang et al., 2024) 30.75 29.50 11.60 4.33 - 5.45
SafePatching (Ours) 7.25 7.50 3.33 2.33 40.42 6.14
Table 1: The overall results on the safety, over-safety and utility benchmarks with LLaMA-2-7B-Chat. The evaluation metrics for safety is the average ASR of four jailbreak methods. Over-safety is evaluated in terms of refusal rate. We report the average results on all utility datasets except for MT-bench (with GPT-4o ratings on a scale of 1 to 10). Results highlighted in green indicate an improvement or performance comparable to the original backbone, while those highlighted in red signal a decline in performance relative to the original backbone.